OpenStudy (slayer8001):

A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix? Make sure to show your work. And use the V.E.S.T. formula

OpenStudy (anonymous):

1-7y56

OpenStudy (slayer8001):

is that all the work?

OpenStudy (slayer8001):

@blasterjakz

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its completely wrong I think

OpenStudy (slayer8001):

I need help fast

OpenStudy (anonymous):

what is the vest formula?

OpenStudy (anonymous):

texas plx help

OpenStudy (texaschic101):

I don't even know the vest formula

OpenStudy (slayer8001):

variable equation solve it and test it

OpenStudy (anonymous):

Let x be the pounds of cashews and y be the pounds of peanuts. Then \(x + y = 10\), since there are 10 pounds total. 5.60x is the amount for cashews and 2.30y is the amount for peanuts. I THINK 3.29(x + y) is the mixture. So the equation is \(5.60x+2.30y=3.29(x+y)\)

OpenStudy (anonymous):

and from there you can use the method of your choice to solve the system

OpenStudy (texaschic101):

wouldn't it just be.. x + y = 10 5.60x + 2.30y = 3.39

OpenStudy (texaschic101):

oops....3.29

OpenStudy (anonymous):

I'm really not sure about the second equation. I included the (x+y) because the 3.29 is per pound

OpenStudy (texaschic101):

hmmm...your right, it is per pound....not too sure about this one

OpenStudy (texaschic101):

@surjithayer ...help

OpenStudy (texaschic101):

I know mine is not right if it is 3.29 per pound.....yours looks right..not 100% sure

OpenStudy (anonymous):

chic

OpenStudy (anonymous):

the wording on this one is tricky

OpenStudy (texaschic101):

yes it is

OpenStudy (anonymous):

very

OpenStudy (texaschic101):

all I can say is tag some ppl

OpenStudy (texaschic101):

@Loser66 ...help

OpenStudy (anonymous):

@master_cheif_117

OpenStudy (texaschic101):

@ganeshie8 ...help

OpenStudy (anonymous):

@kiamousekia

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

frankly speaking i don't know v.e.s.t formula.

OpenStudy (anonymous):

if no loss or profit then equations given by peachpi are correct.